what is the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4.
Answer by pablo9619(2)

(-3,1) parallel to y=-x-4
parallel lines have the same slope
y=mx+b m= slope b= y-intercept
1=(-1)(-3)+b
1=3+b
-3 -3
-2=b
y=-x-2